# A Multimodal Benchmark for Video Fake News Detection

### v1

@@ 4.3 Task definition
According to practical detection process, evaluation tasks can be categorized into three types: 1) perception FND, which requires model to accurately capture fine-grained characteristics in videos. 2) understanding FND, which needs to be solved by observing the global contents in the video. 3) reasoning, which necessitates the joint utilization of general knowledge and information within the video. In this chapter, we present several key task definitions, while the remaining tasks are introduced in the Appendix B.2.
@@ 4.3.1 Perception
1) Key Elements Perception, KEP. The key elements perception task requires MLLMs identify all elements that are crucial for identifying the veracity of the news video. These elements include on-site footage, the task is formatted as open-ended question, all questions are annotated by the process in Figure 4
2) Distinguish Creator-added Content and Original Shooting Footage, DCS. The distinguish task involves recognizing which elements in the video are author-added content and which correspond to the originally captured footage. The question is also annotated by pipeline in the Figure 4. And the format is open-ended generation with a specified format.
3) Creator-added Text Color Perception, CCP. In this task, MLLMs are required to perceive the font color of the provided creator-added text in the input key frame. According to our statistics, the value range of font colors in the video is limited; thus, this task involves fixed-format output with answers constrained within a finite scope. We employ accuracy as the evaluation metric.
4) Creator-added Text 2D-position Perception, CPP. The format of this question is defined as a single-choice type. The input includes 30 position exact descriptions and MLLMs need to recognize and select position of the text within screen based on the key video frame and target text. The performance is measured by accuracy. We create 6.3k Question-Answer pair by sampling key frames and human annotations.
@@ 4.3.2 Reasoning
1) Final Detection Reasoning, FDR. In this task, MLLMs need to utilized the multi-modal information and general knowledge to verify the veracity of videos. This task adopts an open-ended format while adhering to a fixed output structure: the reasoning process should be presented within the <think></think>tags, and the final binary classification result should be outputted within the <result></result> tags. We employ accuracy, entity hit rate, rationale hit rate, knowledge hit rate, theme relevance, factual consistency to evaluate the final detection reasoning result.
Figure 4: Pipeline of dataset construction
Table 2: Evaluation Result of MLLMs on Perception and Understanding in Fake News Detection
Model Perception Understanding
@@
Qwen2.5-VL-72B 59.61 60.26 39.15 50.36 0.2346 2.03 77.27 44.86 4.04 4.08 2.63
Qwen2.5-VL-32B 54.30 57.03 31.71 35.73 0.2218 2.16 70.47 38.67 3.62 3.66 2.27
Qwen2.5-VL-7B 52.86 40.86 28.33 24.82 0.1589 2.71 19.38 30.84 3.37 3.51 1.87
@@
Model Final Detection Reasoning Accuracy
Avg ENHR Avg RHR Avg KHR
Gemini1.5-Flash 62.67 79.57 47.11 78.61
GPT-4o-mini 31.74 41.26 24.75 71.05
Qwen2.5-VL-72B 46.62 68.76 37.59 77.00
Qwen2.5-VL-32B 30.14 61.36 29.85 74.90
Qwen2.5-VL-7B 25.35 40.47 22.73 70.12
InternVL3-78B 25.32 33.09 23.61 75.96
InternVL3-38B 21.17 30.05 21.51 73.43
InternVL3-8B 18.73 26.73 18.64 70.00
Table 3: Final Detection Reasoning results of different models.
5 Experiments
5.1 Setting
@@
We conduct comprehensive experiments on 10 tasks with 6 MLLMs in our benchmark, encompassing both open-source and proprietary models. For image MLLMs, which are primarily fine-tuned using image-related data, we leverage their inherent image process capacities to utilize sampled frames from original news video. For video MLLMs, fine-tuned with video-related instructions, we utilize a uniform sampling method or models’ predefined sampling strategy for video processing. We employ models’ official implementations or available APIs for evaluation. We assess both backbone and fine-tuned models with zero-shot manner. More evaluation-related information can be accessed in the Appendix B.3.
5.2 Main Results
The overall evaluation results for all open-source and proprietary MLLMs in the test set are shown in Table 2 and 3, from the results, we have derived several conclusions for the main results and attach them in the Appendix B.4.
5.3 Discussions
We evaluated the performance of MLLMs across various capabilities on the MVFND task, utilizing both process-based and result-based tasks within the benchmark. Through further analysis of the evaluation process, we identified the following insights and discussed their underlying intrinsic causes.
Insight 1. For MVFND task, the performance bottlenecks of the model vary across different types of news videos. As presented in Figure 5, reasoning relies on different information types across distinct news video categories. It can be observed that knowledge-based news videos rely more heavily on knowledge during the reasoning process, while exhibiting low utilization of entities in the video (e.g., screen text and footage). In contrast, the event-based news videos are less sensitive to external knowledge, and they utilize far more entities derived from the video contents.
Insight 2. For MVFND task, dense frame sampling is not requisite for video content perception. Video frame sampling strategies should take into account both video duration and model input load. As shown in Figure 6, we take the performance of InternVL3-78B as an example. For each video duration group, there exists an optimal number of sampling frames. And this optimal number tends to increase as the video duration extends. But dense frame sampling does not guarantee a continuous improvement in accuracy; instead, it may even compromise the model performance. Dynamic frame sampling strategy yields higher accuracy, especially for shorter news videos.
Insight 3. For MVFND task, better performance can only be achieved with sufficient key elements in the video and the model’s temporal localization capability. As demonstrated in Figure 7, when the number of key elements is one, all models achieve relatively favorable performance. But as the number of key elements in the video increases, only some more advanced models exhibit improved performance.
@@ 5.3 Discussions
Our work has explored the utilization of MLLMs for VFND, however, there are also some limitations Firstly, Fake news instances of the same type may exhibit similar deceptive characteristics, as the nature of the facts they aim to obscure is comparable. This aspect, however, remains underexplored in our current work. Secondly, While we have made every effort to maintain a balanced distribution of news video types within the dataset, there remains room for improvement. type balance should be incorporated as a key evaluation metric for dataset construction in the future works.
Appendix A Empirical Analysis
A.1 Data process
Figure 10: Creator-added Content Extraction Process
Figure 11: Original Shooting Dataset Extraction Process
For empirical analysis, we also divide the data into creator-added contents and original shooting footage. Because creator-added contents carry more author’s intents, so it has less credibility. In order to census more fine-grained data, we construct two data process pipelines to collect creator-added contents and original shooting scenes in details. As shown in Figure 12, The content in the yellow box is the text added by the author, while the content in the red box is the text present in the original footage.
Figure 12: Text region in video screen
Figure 13: Individual with identity
Most of author-added elements in news video are text, the creators utilize specialized software tools to insert textual elements as complementary information. For example, adding captions to explain event progression and introduce key figures, sometimes authoritative medias may exhibit distinctive graphical identifiers. Apart from discrepancies in the Hue, 2D-position of creator-added contents(CAC), there are also a few characteristics of video news have demonstrate discrepancy. We add other analysis towards to fake and real video news in the following parts. As the Figure 10 shown,For creator-added contents process pipeline, we uniformly sampled 16 frames per video using FFmpeg, maintaining temporally equidistant intervals between extracted frames.After that, we leverage the OCR function of Paddle to extract text regions of each frame. The outputs of Paddle include bounding box and screenshots. However there are also a few text regions extracted from original shooting footage, such as bus stop marker. So we manually removed these texts and get all creator-added text regions. Then, we design prompt templates and employ qwen2.5vl-72b to recognize and generate semantic description of text content, color, 2D-position and aspect ratios. Finally, experts refine the descriptions generated by MLLMs. Figure 11 demonstrates the overall process of extraction, generation and refinement.
Figure 14: Distribution of area
in fake/real news video
Figure 15: Distribution of aspect ratio in fake/real news video
Textual information is usually placed on video screen by authors to deliver key descriptions about the news. However, in order to minimize text occlusion with the original shooting scenes and maximize the transmission efficiency of both textual and visual information, the text region should be designed empirically with above considerations. To analyze the outline discrepancy of text regions in fake/real news video, we census the area size and aspect ratio of them. As shown in Figure 14, real videos are more likely to utilize a moderate region size, while fake videos have a larger proportion in both relatively large and small parts. Such comparison result aligns with practical observations, where larger text regions are often employed by false information creators to divert viewers’ attention toward misleading text content or to occlude authentic details in the original captured scene. Smaller text regions typically emerge when fake news producers have limited relevant news material of the news. The Figure 15 demonstrates that aspect ratio of text regions in news videos serves as a distinct delimiter for differentiating between authentic and fabricated content. Specifically, text regions’ aspect ratio in fake news videos tend to be smaller than 15, whereas those in real news videos typically have an aspect ratio larger than 15.
A.1.1 Color Distribution
Figure 16: Saturation distribution in creator-added content of fake/real news video
Figure 17: Value distribution in creator-added content of fake/real news video
In color theory, hue, saturation, and value (HSV) are the fundamental attributes of each color. Color is typically employed to emphasize key information in video news. Saturation describes the intensity or purity of a color, where varying its value yields distinct visual effects even within the same hue. Value (or brightness) represents the perceived luminance, with higher values indicating a greater proportion of white light in the color. The combined numerical representation of these three dimensions provides a unique discriminative signature for each color. We utilize these factors to describe creator-added text in pixel level. We have already show the difference of hue in fake and real video news with the Figure 1, and we will continue to demonstrate the discrepancy of saturation and value in Figure 16 and Figure 17. As shown in Figure 16, fake video news have lower saturation compared with fake news video when saturation lower than 0.35. And the relationship of fake and real video news has reversed when saturation larger than 0.35. However, the difference between real and fake video news is not clear.
Figure 18: Arousal and Dominance scores of materials distribution
A.1.2 Emotion of audience while watching the video
Fake news intend to change audience’s emotion, as a result, the cognition may change along with the corresponding emotion. So we utilize the MLLMs to emulate audience watching video news, assessing audience’s arousal and dominance level with scres. As demonstrated in Figure 18, real video news arouse high arousal and low dominance, which proves real news can utilize moderate emotional level to stimulate higher arousal, because real news have more valuable information that have clear emotional influence. However, fake news give audience higher dominance that reprensents human’ s feeling over sense of control, People are more willing to make decision when dominance is high. So we discover fake news video creates an illusion of high control for the audience. so much so that the audience believes the content in the video is true.
A.2 Original shooting footage
=Original shooting footage usually is the first hand description of event, especially sudden incident. Most of the footage is dynamic, so we utilize Gemini to process nascent videos. For key event perception, first we leverage MLLMs to process video with prompt template, after that experts manually optimize the outputs. For key event grounding, we employ MLLMs to locate key events with outputs of the perception, then experts adjust the time ranges of key events generated by MLLMs. For human identity perception, we use MLLMs to retrieve figures related to news event. Then experts manually revise the results.
A.2.1 Subject Identity Distribution
Figure 19: Distribution of subject identity in
fake(external)/real(internal) news video.
News videos frequently feature human subjects to convey key event information through specific visual formats, such as face-to-face interviews and press conference footage. These individuals include involved parties, victims, law enforcement personnel, self-media creators, and others. However, the reliability of the information provided by these identities varies significantly Pereira et al. (2023). To investigate this, we analyzed the proportional distribution of different identities across the entire duration of the videos. Detailed results are presented in the Appendix. We find the perpetrator, victim and office personnel appear more frequently in real news video. However, larger proportion of fake news videos contain no people and the we-media appears frequently in the fake news. For the distribution of subject identity shown in Figure 19, we find the perpetrator, victim and office personnel appear more frequently in real news video. Especially law enforcement officers who appear in real news videos more than twice as frequently as in fake news videos. We attribute this significant disparity to real news producers can access law enforcement personnel involved in reported events. Moreover, the inclusion of such content is an essential procedure to alleviate audience skepticism. Conversely, fake news producers actively avoid this element. However, larger proportion of fake news videos contain no people and the we-media appears frequently in the fake news. The result align with our practical observation, because fake news generally has less relevant subjects to enforce the veracity of the video news.
Figure 20: Distribution of category-test-split
Figure 21: Number of shooting angles in the fake and real video news
A.2.2 Relevant shooting angles
Finding In terms of shooting angles, fake news tends to exhibit fewer camera shot transitions, whereas real news shows the opposite pattern.
Analysis To provide a comprehensive and objective depiction of news events, creators often employ multiple shooting angles within their videos. This approach not only enables audiences to gain a thorough understanding of the event’s occurrence and progression but also enhances the credibility and persuasiveness of news reports Bock et al. (2023). We attribute the discrepancy to the reasons, 1) fake news producers typically have no enough shooting materials about the events. 2) less shooting footage prevent audience from understanding the fact. 3) producers of fake news lack the professional skills required for camera shot transitions in the process of news video production.
Appendix B Experiments
B.1 Data Curation
Figure 22: Evaluation process of MLLMs’ outputs
B.1.1 Details of Videos Category

### v2

@@ 3.1 Task definition
Following practical detection workflows, we design tasks targeting different stages of the verification process. Perception tasks target the stage of visual evidence extraction from both creator-added content(CAC) and original shooting footage(OSF). Understanding tasks target the global video news comprehension. Reasoning tasks target the stage where MLLMs synthesize evidence to produce verification judgments. In this section, we present formal task definitions and discuss their necessity in the VFND. All tasks are annotated by the process in Figure 3, and concrete examples of these tasks are provided in Appendix 10.
Figure 3: POVFND Dataset Construction.
@@ 3.1.1 Perception
Key Elements Perception (KEP). The task requires MLLMs to identify all elements in original shooting footage that are crucial for VFND. These elements include on-site footage, interview segments, and other components. These elements offer primary source information about the events.
Distinguish Creator-added Content and Original Shooting Footage (DCS). The task involves recognizing creator-added content and original shooting footage in video news. The former is more susceptible to manipulation and carries stronger authorial intent, whereas the latter provides relatively unaltered information. However, MLLMs usually treat both types as equally credible, as they are pretrained on semantic features rather than provenance distinctions, as illustrated in Figure 4.
Creator-added Text Color Perception (CCP). In this task, MLLMs are required to perceive the font color of text in videos. Text color constitutes a crucial component of news, as it reflects the creator’s professional background and influences comprehension Zhou et al. (2022). However, previous experiments reveal that Clip-based MLLMs struggle to identify color of visual content, as the error illustrated in Figure 5.
Creator-added Text 2D-position Perception (CPP). Similar to text color, text position reflects the creation style of news content, yet MLLMs face limitations in recognizing spatial locations of text.
Shooting Angles Counting(SAC). More camera angles enable multi-perspective record of news events, particularly in accident scene footage. Thus, camera angles constitute a crucial attribute of OSF. This task requires MLLMs to identify angle transition and count the number of distinct viewpoints.
Human Identity Recognition (HIR). Given a video, MLLMs were required to retrieve and identify all identities that have appeared in the video.
Key Elements Temporal Grounding (KEG). By leveraging the temporal ranges of key elements, MLLMs conduct more precise visual information extraction and verification. Consequently, KEG enhances MLLMs’ video understanding accuracy, thereby improving overall VFND performance.
3.1.2 Understanding
News Video Type Understanding (NTU). Different news types rely on distinct evidence for verification, as illustrated in the Figure 19. Therefore, determining video type before applying type-specific detection strategies can effectively improve both VFND efficiency and accuracy.
News Video Theme Understanding (NEU). This task requires MLLMs to obtain a global comprehension of video news and extract thematic information (What, When, Where, Who, How, etc.). This information guides MLLMs to focus on elements relevant to the key content.
@@ 3.1.3 Reasoning
Final Detection Reasoning (FDR). This task requires MLLMs to synthesize multi-modal entities extracted through perception and understanding capabilities, incorporate general knowledge, and apply logical rationales to reason video veracity.
To comprehensively evaluate MLLMs’ capabilities, we employ task-specific evaluation strategies tailored to output formats, as shown in Table 2. For structured outputs, we use rule-based exact matching metrics(e.g.,IoU). For instance, in the CCP task, we assess performance via perception accuracy. For tasks requiring holistic comprehension of global video semantics, such as news theme understanding, we adopt semantic-based metrics including factual consistency, relevance, and completeness, evaluated through GPT-4o as an automated judge.
Figure 4: When identifying video veracity, MLLMs uncritically accept the text added by creators, which leads to erroneous conclusions.
3.2 Dataset Construction and Statistics
Data Collection. We collect videos from FakeSV, which contains real-world news videos from popular social media platforms across diverse domains. This dataset offers two key advantages: (1) labels are cross-validated against official fact-checking platforms, ensuring reliability; (2) videos retain authentic social media characteristics, enabling our benchmark to capture realistic fake news patterns and align with practical VFND. To ensure unbiased evaluation, we construct a balanced test set with 753 video pairs across 6 distinct news categories, knowledge-based news account for 13.4%.
Data Annotation. For data annotation, as depicted in the Figure 3, to reduce bias and hallucination, we utilize Gemini2.5-flash with strong video abilities to generate and refine visual tasks. Then, we use O1 to evaluate and filter the reasoning outputs of MLLMs. The O1 model evaluates these reasoning outputs for factual consistency and logical rationality, assigning scores from 0 to 5. Outputs scoring below 4 in either metric are reprocessed by the Gemini. Finally, perception generations are verified by 20 news creation experts after 10 hours annotation training; understanding and reasoning outputs are reviewed and revised by 10 detection experts (with 5+ years of video news censor experience in online public opinion regulatory authority) and 5 domain experts(with 5+ years of specialized field research experience). This iterative process continues until all task data is processed.
Figure 5: CLIP model misclassifies text color in video; green is given higher similarity.
Data Statistics. POVFNDB comprises 3,624 videos across 6 news categories, split into 753 test videos and 2,871 training videos. Each video is accompanied by visual information, semantic descriptions, and reasoning evidence. As shown in Figure 6, the test split contains an average of 4.40 key elements per video, totaling 2,875 key elements, such as on-site footage. The duration of key elements and the number of key shots exhibit diverse distributions. The dataset maintains balanced representation across 6 news categories.
Figure 6: POVFND Dataset Statistics.
3.3 Evaluation Metrics
To evaluate MLLM performance across the 10 tasks, we define task-specific metrics tailored to output formats and evaluation objectives. In this section, we present the mapping between metrics and tasks, with detailed calculation procedures provided in Appendix G.
Avg. EHR (Element Hit Rate): Measures the element hit rate in the KEP task, assessing MLLMs’ ability to identify critical visual information in news videos (e.g., on-scene footage).
Avg. ENHR (Entity Hit Rate): Measures the entity hit rate in the FDR task, assessing MLLMs’ ability to cite video entities during reasoning (e.g., on-screen text for locations and identities).
Avg. KHR (Knowledge Hit Rate): Measures the knowledge hit rate in the FDR task, assessing MLLMs’ ability to leverage external knowledge during reasoning (e.g., physics, economics).
Avg. RHR (Rationale Hit Rate): This metric measures the rationale hit rate in the FDR, evaluating MLLMs’ ability to apply reasoning logic during the process, such as contradictions between video content and facts that reduce credibility.
Avg. AD (Absolute Distance): Measures the absolute distance between predicted and ground truth shooting angle counts in the SAC task. Lower values indicate better performance.
Avg. IoU (Intersection over Union): Measures temporal grounding accuracy in the KEG by evaluating MLLMs’ ability to localize multiple key elements per video. Higher values indicate better localization performance.
4 Experiments
  Model Task Perception     Understanding
@@ 4 Experiments
Qwen2.5-VL-72B-Instruct 59.61 60.26 39.15 50.36 0.2346 2.03 77.27     44.86 4.04 4.08 2.63
Qwen2.5-VL-32B-Instruct 54.30 57.03 31.71 35.73 0.2218 2.16 70.47     38.67 3.62 3.66 2.27
Qwen2.5-VL-7B-Instruct 52.86 40.86 28.33 24.82 0.1589 2.71 19.38     30.84 3.37 3.51 1.87
@@
Table 3: Evaluation on Perception and Understanding. Best results are in bold with darker background, second best with lighter background.
  Model Task     Reasoning
    ENHR RHR KHR ACC
Gemini2.5-Flash     62.67 79.57 47.11 77.29
GPT-4o-mini     31.74 41.26 24.75 69.72
Qwen2.5-VL-72B-Instruct     46.62 68.76 37.59 75.70
Qwen2.5-VL-32B-Instruct     30.14 61.36 29.85 73.17
Qwen2.5-VL-7B-Instruct     25.35 40.47 22.73 68.79
InternVL3-78B     25.32 33.09 23.61 74.50
InternVL3-38B     21.17 30.05 21.51 71.71
InternVL3-8B     18.73 26.73 18.64 68.26
Table 4: Evaluation on Final Detection Reasoning.
MethodTask     Qwen2.5VL-7b-Instruct
    ACC M-F1 ENHR RHR KHR
Zero-Shot     68.79 67.41 25.35 40.47 22.73
CoT POVFND     70.14 68.54 28.37 41.57 24.39
CAC-POVFND     71.58 70.64 32.15 45.65 26.71
OSF-POVFND     73.87 72.38 33.17 49.26 29.71
ALL-POVFND     72.64 71.59 40.13 48.31 28.10
Instruct-Tuning     81.14 80.26 74.58 83.15 51.37
Table 5: Evaluation results of different CoT variants on VFND using Qwen2.5VL-7b-Instruct.
Utilizing POVFNDB, we evaluate a diverse set of MLLMs, including both video-based and image-based models from open-source and proprietary source. Overall evaluation results are presented in Table 3 and 4. Beyond the process-oriented tasks evaluation, we select rationales in MLLMs’ reasoning outputs, conducting in-depth analysis within video visual features to verify validity of these rationales. Based on these findings, we propose POVFND-CoT (Table 9) to assess performance gains from feature integration on detection accuracy. We further fine-tune Qwen2.5VL-7B-Instruct on the reasoning outputs generated by POVFND-CoT. The fine-tuned model achieves sota performance across 4 reasoning evaluation metrics.
4.1 Experimental Settings
@@
We conduct comprehensive experiments on 10 tasks with 8 MLLMs in our benchmark, encompassing both open-source and proprietary models. For image MLLMs, we leverage their inherent image process capacities to utilize evenly sampled 16 frames from each news video. For video MLLMs, we utilize a uniform sampling method or models’ predefined sampling strategy for video processing. We employ models’ official implementations or available APIs for evaluation. We assess both backbone and fine-tuned models with zero-shot manner.
4.2 Results Analysis
News Video [Uncaptioned image]
Title Belt and Road Land Freight Train from Yiwu, Zhejiang to London, UK
Original Footage & Creator Content Original footage: Brands on the train. …train moving through rural landscape…. Creator-added content:Belt and Road land transport train from Yiwu to London”.
Creator-added Text Content: Belt and Road … from Yiwu… to London,…; Color: white; Position: center and bottom center.
Key Elements Perception Branding on train engines: BNSF, HYUNDAI and MAERSK logos. Environmental context: Landscape, vegetation, railway infrastructure…
Key Elements Grounding & Analysis 20s-30s (BNSF):“BNSF” logo. BNSF = Burlington Northern Santa Fe, a major North American railroaddoes not operate on Eurasian land bridgeGeneral landscape: … railway infrastructure aligns with North American standards
Identity Recognition No specific individuals (interviewees, officials, reporters) appear. Purely train and landscape footage.
Shooting Angles Two angles capturing train passing by, likely from fixed position or drone.
POVFND-CoT Veracity Reasoning <think>…the "BNSF" logo belongs to Burlington Northern Santa Fe, a major North American railroad operating exclusively in the United States and Canada, not on China-Europe routes … HYUNDAI and MAERSK markings, neither of which operates direct rail services from Yiwu to London. …, railway signals, and silos characteristic of the North American Midwest, not the Eurasian…… The text’s use of "supposedly" is highly unusual for factual news reporting, indicating unverified sources. The non-standard font and intrusive text layout occupying half the screen further deviate from conventional news presentation standards. The fundamental discrepancy between the BNSF North American footage and the claimed China-Europe narrative, combined with stylistic red flags, clearly identifies this as fake news through content misappropriation</think><result>Fake</result>
Ground Truth Fake
Table 6: Case study demonstrating POVFND-CoT’s step-by-step reasoning process.
MLLMs exhibit poor performance on creator-added text tasks (CCP and CPP), with the best model achieving below 60% accuracy on both tasks. We attribute this to MLLM training paradigms that prioritize semantic understanding while neglecting spatial positioning and visual attributes (e.g., font color). This reveals a critical gap in existing MLLMs for news video perception, where text-rich visual content is prevalent. Our results suggest that specialized training targeting text perception abilities is essential for effective MLLM-based VFND.
Temporal grounding across multiple key elements challenges model detection performance. Only Gemini-2.5-Flash exceeded the standard threshold of 0.3, indicating that dense temporal grounding in news videos remains challenging for off-the-shelf MLLMs. Figure 22 demonstrates that all models achieve reasonable performance with a single element. However, as the number of key elements increases, only models (e.g., Gemini2.5-Flash) with stronger localization abilities show performance gains. Notably, we observe a positive correlation between multi-element grounding capability and VFND accuracy.
Video-based MLLMs demonstrate better dynamic feature capturing capabilities. For instance, Qwen2.5-VL-7B achieves comparable temporal grounding performance to GPT-4o-mini while exhibiting stronger SAC capabilities, despite its smaller size. We attribute the advantage to Qwen2.5-VL’s native video processing architecture and dynamic FPS sampling mechanism, which enable more effective temporal modeling. These findings suggest that image-based processing paradigms limit MLLMs’ visual abilities and VFND performance.
Hit rate metrics provide evidence-based reasoning verification. As demonstrated in Figure 4, hit rates exhibit strong positive correlation with detection accuracy, with improvements in hit rates consistently accompanied by accuracy gains. These patterns align with real-world verification practices, indicating that our benchmark closely reflects practical detection workflows.
OSF features provide larger improvements than CAC. As shown in Table 5, we attribute this to two factors. First, OSF captures dynamic visual elements with richer contextual information, while CAC is limited to semantic with a constrained feature space. Second, OSF originates from authentic footage and is less susceptible to manipulation, whereas CAC represents author-added content with lower credibility. Consequently, models rely more heavily on OSF, yielding superior detection performance. This suggests that MLLM-based VFND should prioritize exploring discriminative patterns in OSF when conduct authentic VFND.
4.3 Further Analysis
Based on evaluation results and rationale validation, we propose POVFND-CoT (Table 9), which tailors reasoning paths based on news type: knowledge-oriented, content-oriented, or hybrid. This CoT approach enables MLLMs to cite external knowledge, perceive target visual features, and apply detection rationales to assess veracity. To examine how different video features contribute to detection performance, we develop CoT variants combining POVFND with CAC and OSF: CAC-POVFND-CoT : POVFND integrates creator-added content analysis (color, position, text), details as shown in Appendix Table 11. OSF-POVFND-CoT : POVFND incorporates original shooting footage analysis (key elements, shooting angles, human identities), details as shown in Appendix Table 12. ALL-POVFND-CoT : POVFND integrates both CAC and OSF components, details as shown in Appendix Table 10.
As shown in Table 5, we find MLLMs struggle to integrate OSF and CAC., ALL-POVFND-CoT underperforms OSF-POVFND-CoT despite incorporating CAC information. We attribute this to MLLMs’ inability to properly weight each feature’s contribution, reflecting VFND’s domain-specific reasoning abilities.
To address this, we use ALL-POVFND-CoT to guide Gemini2.5-Flash in generating reasoning outputs. This enables MLLMs to utilize extensive visual information and perform end-to-end detection, incorporating retrieval triggers and evidence-based rationales. The outputs reveal how MLLMs weigh various evidence to derive conclusions. We fine-tune Qwen2.5-VL-7B-Instruct on these training samples (parameters in Table 8), with the fine-tuned model achieving state-of-the-art performance in test-split. This demonstrates that effective VFND requires specialized reasoning strategies developed through domain-specific fine-tuning rather than general capabilities alone.
4.4 Case Study
Does the MLLM distinguish between creator-added text and text in the original footage ? As shown in Table 6, the model identifies text on the train cars, such as ’BNSF’, as well as the creator-added text ’Belt and Road land transport train from Yiwu to London’. The MLLM utilizes these two distinct text types for different analytical purposes. For creator-added text, it examines the location, font color, and semantic content. For text appearing on the train, the model uses it to retrieve information for fact-checking.
Does the MLLM effectively leverage external knowledge to assist VFND ? The model utilizes the captured logo ’BNSF’ on the train to retrieve the fact that ’BNSF is a major North American railroad operating exclusively in the United States and Canada, not on China-Europe routes’. Additionally, the MLLM analyzes the environmental context in the footage and identifies characteristics typical of the North American Midwest rather than Eurasia. These factual discrepancies with the news content significantly undermine the credibility of the news.
Does the MLLM correctly identify the news type as knowledge and content oriented ? The model’s output incorporates both external knowledge verification and video content analysis, indicating that the MLLM recognizes the news type during reasoning and adapts its subsequent reasoning steps.
5 Conclusion
In this work, we introduce POVFNDB, a process-oriented benchmark for evaluating MLLM capabilities in VFND through comprehensive video feature descriptions and process-centric tasks. Our extensive evaluation reveals significant limitations in existing MLLMs’ perception, understanding, and reasoning abilities for fake news detection. We identify key strengths and weaknesses of current MLLMs and establish that effective VFND necessitates domain-specific reasoning strategies rather than relying solely on general-purpose capabilities.
@@ 5 Conclusion
Our work has explored the utilization of MLLMs for VFND, however, there are also some limitations Firstly, Fake news instances of the same type may exhibit similar deceptive characteristics, as the nature of the facts they aim to obscure is comparable. This aspect, however, remains underexplored in our current work. Secondly, the application of external knowledge requires careful verification of its authenticity, which typically necessitates cross-validation from multiple sources. Further exploration of these approaches remains necessary.
Appendix A Data Curation
A.1 Details of Videos Category